# Collective Variable-Guided Engineering of the Free-Energy Surface of a Small Peptide

### v1

@@ Abstract
Engineering the free-energy surfaces (FES) of proteins and peptides is central to controlling conformational ensembles and their responses to perturbations. However, predicting how chemical modifications such as point mutations reshape the FES and shift conformational equilibria remains challenging, particularly in data-scarce settings. Building on the Collective Variables for Free Energy Surface Tailoring (CV-FEST) framework, we develop a computational approach that leverages short, unbiased molecular dynamics trajectories to guide mutation analysis. Using the ten-residue $\beta$-hairpin CLN025 and a systematic library of its single-point mutants, we apply Harmonic Linear Discriminant Analysis (HLDA) to extract collective variables from the conformational data. We find that the HLDA eigenvector learned solely from short wild-type trajectories provides residue-level insight into the propensity of mutations at specific positions to thermodynamically stabilize or destabilize the folded state. Extending this analysis, we show that shifts in the leading HLDA eigenvalue across mutants—a measure of changes in separability between the conformational ensembles along the HLDA coordinate—correlate strongly with mutation-induced changes in the free-energy difference between states, as reflected in melting temperatures. Benchmarked against Replica Exchange Molecular Dynamics simulations, these findings suggest a promising and computationally affordable route toward guiding the engineering of biomolecular free-energy landscapes.
@@ Introduction
To address these limitations, we explore in this work a novel approach to the engineering of protein and peptide FES, with particular emphasis on data-scarce regimes in which large deep-learning (DL) models are often less effective. Our approach builds on the recently developed Collective Variables for Free Energy Surface Tailoring (CV-FEST) framework Mendels and De Pablo (2022); Mendels et al. (2023), which learns collective variables (CVs) from MD data to capture slow, relevant conformational transitions. CV-FEST is specifically suited to low-throughput settings, where only a limited number of FES calculations are feasible.
@@ Acknowledgements.
Data Availability
Data used in this work is available upon request.

### v2

@@ Abstract
Engineering the free-energy surfaces (FES) of proteins and peptides is central to controlling conformational ensembles and their responses to perturbations. However, predicting how chemical modifications such as point mutations reshape the FES and shift conformational equilibria remains challenging, particularly in data-scarce settings. Building on the Collective Variables for Free Energy Surface Tailoring (CV-FEST) framework, we develop a computational approach that leverages short, unbiased molecular dynamics trajectories to guide mutation analysis. Using the ten-residue $\beta$-hairpin CLN025 and a systematic library of its single-point mutants, we apply Harmonic Linear Discriminant Analysis (HLDA) to extract collective variables from the conformational data. We find that the HLDA eigenvector learned solely from short wild-type trajectories provides residue-level insight into the propensity of mutations at specific positions to thermodynamically stabilize or destabilize the folded state. Extending this analysis, we show that shifts in the leading HLDA eigenvalue across mutants, a measure of changes in separability between the conformational ensembles along the HLDA coordinate, correlate strongly with mutation-induced changes in the free-energy difference between states, as reflected in melting temperatures. Benchmarked against Replica Exchange Molecular Dynamics simulations, these findings suggest a promising and computationally affordable route toward guiding the engineering of biomolecular free-energy landscapes.
@@ Introduction
To address these limitations, we explore in this work a novel approach to the engineering of protein and peptide FES, with particular emphasis on data-scarce regimesSalman et al. (2025) in which large deep-learning (DL) models are often less effective. Our approach builds on the recently developed Collective Variables for Free Energy Surface Tailoring (CV-FEST) framework Mendels and De Pablo (2022); Mendels et al. (2023), which learns collective variables (CVs) from MD data to capture slow, relevant conformational transitions. CV-FEST is specifically suited to low-throughput settings, where only a limited number of FES calculations are feasible.
@@ Acknowledgements.
Data and Software Availability
All scripts and input files required to reproduce the analyses in this study are publicly available at https://github.com/Muralika-M/CV-guided-FES-analysis. The repository contains Python scripts for HLDA analysis, threshold scanning, correlation analysis, and figure generation, together with representative example datasets and the GROMACS and PLUMED input files used in this work. Molecular dynamics trajectories are not included due to their size but can be provided by the authors upon reasonable request.